OpenStudy (anonymous):

|-15| + 5^2 = 70y - |-33| - 69y -7??

OpenStudy (hoblos):

|-15| + 5^2 = 70y - |-33| - 69y 15 +25 =y -33 y=73

Directrix (directrix):

73 is what I get.